### v1

@@ 2.1 Network architecture
While neural networks have been successfully used in stock return forecasting (Gu et al., 2020), quantile forecasting is challenging because we need to capture both cross-sectional variation across stocks and market-wide variation over time. One way to address this challenge is to use cross-sectional standardisation of returns to remove the time series noise. This is possible from an asset pricing perspective, where the goal is to rank individual stocks cross-sectionally to create long-short portfolios. However, since our goal is to forecast quantiles of future raw (non-standardised) stock returns, we propose a two-stage neural network architecture that exploits the predictability of standardised returns in the first stage and generates quantile forecasts of true (non-standardised) stock returns in the second stage. In this way, we allow the neural networks to exploit the information in standardised returns and to adjust for expected market volatility.
@@
where $\overline{\sigma}_{t}=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\sigma_{i,{t}}$ is a cross-sectional average of stock-level volatilities, $\sigma_{i,t}$ is a stock-level volatility estimate of stock $i$ at period $t$, and $N_{t}$ is the number of stocks in the panel at period $t$. Ideally, we would use $\overline{\sigma}_{t+1}$ instead of $\overline{\sigma}_{t+1}$ to rescale the standardised quantile forecasts back to the raw quantile forecasts of stock returns, but this is not known at period $t$ when the forecast is made. Therefore, we use the cross-sectional average of stock-level volatilities at period $t$ instead of the future value of $\overline{\sigma}_{t+1}$ to avoid look-ahead bias. $\sigma_{i,t}$ is estimated using an exponentially moving weighted average of the squared returns of stock $i$ with a decay factor of $0.94$ and is one of the input variables to the model. [^5] The 0.94 parameter is widely used in both academia and the financial industry, an example being the RiskMetrics GARCH specification. See Appendix C.1 for a visualisation of $\overline{\sigma}$ over time in the US.
The idea behind standardising individual stock returns using the cross-sectional average of stock-level volatilities rather than the stock volatility itself is to reduce the noise from individual stock volatility estimates. For the same noise-reducing reason, we standardise stock returns only by dividing by the average volatility and and not by subtracting the average return, either stock or market, because average returns are not persistent.
A stock return quantile is then modelled by the two-stage quantile neural network with two sub-networks as
$Q_{r_{i,t+1}}^{\mathcal{T}}$ $=$ $\Big{\{}Q_{r_{i,t+1}}(\tau_{1}|x_{t},z_{t},\overline{\sigma}_{t}),\ldots,Q_{r_{i,t+1}}(\tau_{K}|x_{t},z_{t},\overline{\sigma}_{t})\Big{\}}$ (2)
$=$ $\Bigg{[}\Big{(}\underbrace{g^{(L)}_{W_{S_{1}}^{(L)},b_{S_{1}}^{(L)}}\circ\ldots\circ g^{(1)}_{W_{S_{1}}^{(1)},b_{S_{1}}^{(1)}}\left(x_{t}\right)}_{\text{standardized $\tau$ quantiles sub-network}}\Big{)}\times\overline{\sigma}_{t}\Bigg{]}\times\Big{(}\underbrace{g^{(2)}_{W_{S_{2}}^{(2)},b_{S_{2}}^{(2)}}\circ g^{(1)}_{W_{S_{2}}^{(1)},b_{S_{2}}^{(1)}}\left(z_{t}\right)}_{\text{market volatility sub-network}}\Big{)}$
$=$ [$\underbrace{\Bigg{[}\Big{\{}\underbrace{Q_{\widetilde{r}_{i,t+1}}(\tau_{1}|x_{t}),\ldots,Q_{\widetilde{r}_{i,t+1}}(\tau_{K}|x_{t})}_{\text{Stage I}}\Big{\}}\times\overline{\sigma}_{t}\Bigg{]}\times\widehat{\sigma}_{t}^{M}}_{\text{Stage II}},$
@@
It is important to note that, in contrast to the literature, we consider a multi-output (deep) neural network to characterise the collection of quantiles. Our network has two subnetworks, namely a standardised $\tau$ quantile network and a market volatility network in which $l\in{1,...,L}$ hidden layers transform the input data into a chain using a collection of nonlinear activation functions $g_{S}^{(1)},\ldots,g_{S}^{(L)}$. An activation function, $g^{(\ell)}_{W_{S}^{(\ell)},b_{S}^{(\ell)}}$, is used as a
@@
$\overline{\sigma}_{t}$ $Q^{\mathcal{T}}_{\widetilde{r}_{i,t}}$ (37$\times$1) $\times$$\times$$\times$(128$\times$2)(128$\times$2) Bottleneck (4$\times$1) $x_{t}$ (176$\times$1) $\times$$\times$$\times$ $z_{t}$ (18$\times$1) (8$\times$1) $\widehat{\sigma}_{t}^{M}$ (1$\times$1) $Q^{\mathcal{T}}_{r_{i,t}}$ (37$\times$1) Stage IStage IIStandardized $\tau$-quantiles sub-networkMarket volatility sub-network
Figure 2: Two-stage neural network architecture. Stage I generates cross-sectionally standardised $\tau$-quantile forecasts, $Q^{\mathcal{T}}_{\widetilde{r}}$, as output. Cross-sectional standardisation is performed by dividing individual future stock returns $r_{i,t+1}$ (labels) by the cross-sectional average of the individual stock standard deviation estimates, $\overline{\sigma}_{t+1}$. The standard deviation estimate is an exponential weighted moving average of volatility with a smoothing factor of $0.94$. Step II uses $\overline{\sigma}_{t}$, i.e. based on data from the previous period, to rescale the standardised $\tau$-quantile forecasts generated in Stage I back to the original scale. The rescaling takes place at the multiplication nodes marked with $\times$, representing the multiplication operation between the between the inputs of the node. The market volatility part then generates forecasts of the next period’s market volatility in order to to further refine the $\tau$-quantile forecasts produced as the output of stage II. The forward pass generates a tensor of form $(B,37,2)$, where $B$ is the number of observations in the batch, 37 is the number of predicted quantiles, and 2 corresponds to the standardised and raw versions.
Figure 2 details the architecture of the two-stage quantile neural network model defined above. This is our main model used for quantile forecasting, and we refer to it as the two-stage model throughout the paper. The first stage of the model is used to forecast quantiles of cross-sectionally standardised next month stock returns $Q^{\mathcal{T}}_{\widetilde{r}_{i,t}}$. We use $37$ $\tau$s in the set $\mathcal{T}$ ranging from the extreme left tail of the distribution (0.00005) to the extreme right tail of the distribution (0.99995). The second stage of the model generates quantile forecasts of the next month’s raw stock returns $Q_{r_{i,t}}(\tau)$ for $\tau\in\mathcal{T}$. This stage consists of three steps. First, the standardised quantile forecasts $Q^{\mathcal{T}}_{\widetilde{r}_{i,t}}$ from the first stage and the cross-sectional average of stock-level volatilities $\overline{\sigma}_{t}$ are used as inputs to rescale the standardised quantile forecasts back to the raw stock return scale $Q_{r_{i,t}}(\tau)=Q_{\widetilde{r}_{i,t}}(\tau)\cdot\overline{\sigma}_{t}$. Second, the market volatility subnetwork generates a second scaling factor to reintroduce the time-series market-wide volatility component. This market volatility scaling factor is the same for all stocks in a given period $t$ and is used to refine the quantile forecasts rescaled in the first step. Finally, both the standardised and raw quantile forecasts are clipped at -100% during the forward pass of the neural network.
@@ 2.2 Training with quantile loss function
In order to jointly predict multiple $\tau$-quantiles using a two-stage neural network, $\tau$-quantile losses for individual $\tau$s must be aggregated into a single loss. The two-stage quantile neural network model also requires both standardised and raw stock returns as labels, so the aggregated quantile loss must take this into account. The loss function used during training is defined as a linear combination of the aggregated quantile losses from the first and second stage outputs, as shown in Equation 3. Both coefficients in the linear combination are equal, implying equal weighting of both stages in the loss function. [^6] This specification worked out of the box, and the hyperparameter was therefore not calibrated. In practice, the first stage has a much larger impact on the loss function because it focuses on rescaled raw returns with a scale close to one, while the second stage focuses on raw returns with a scale of about 0.1. The greater implicit focus on the first stage is desirable, as this is where most of the learning takes place, with the second stage only fine-tuning the results of the first stage. The aggregated multi-$\tau$-quantile loss used in the estimation is defined as
@@
where $\widehat{Q}_{r_{i,t}}(\tau)$ is the forecast $\tau$-quantile of the raw return of stock $i$ at period $t$, $\widehat{Q}_{\widetilde{r}_{i,t}}(\tau)$ is the predicted $\tau$-quantile of the standardised return of stock $i$ at period $t$, $B$ is the lot size and $K=37$ is the number of $\tau$ in the $\mathcal{T}$ set. An essential part of the loss is $\rho_{\tau}$, a quantile loss function defined as a piecewise linear function $\rho_{\tau}(\xi)=\mathds{1}_{(\xi\geq 0)}\tau\xi+\mathds{1}_{(\xi<0)}(\tau-1)\xi$ where $\xi=y-\widehat{Q}_{y}(\tau)$ is an error term between the true value and the predicted $\tau$ quantile, and $\mathds{1}_{(.)}$ is an indicator function. A single $\tau$ quantile loss calculated for multiple observations is then calculated as a simple arithmetic mean of the individual $\rho_{\tau}$ losses. The idea behind the quantile loss function is to penalise overestimation of the $\tau$ quantile by $\tau$ times the error if the error is positive and by $(\tau-1)$ times the error if the error is negative. For example, if $\tau=0.5$, the quantile loss function is the absolute value loss function, and overprediction and underprediction are penalised equally. If $\tau=0.05$, underprediction is penalised $19$-times $(0.95/0.05)$ more than overprediction, and minimising such a loss function leads to the estimation of the $0.05$ quantile.
This definition implies an equal weighting of individual $\tau$-quantile losses in the aggregate loss function. In our case, this means that large stocks have the same impact as small stocks and central quantiles have the same impact as tail quantiles.
From a numerical point of view, the quantile loss function is a piecewise linear function that is differentiable everywhere except at zero. Gradient-based optimisation algorithms require the loss function to be differentiable in order to update the model parameters. Automatic differentiation engines, such as PyTorch’s autograd (Paszke et al. (2019)) are able to deal with this problem by defining subgradients for non-differentiable points. A subgradient at a point is any slope of a line that lies below the function at that point and intersects the function at that point. For the quantile loss function, the subgradient at zero can be any value between the left derivative ($\tau-1$) and the right derivative ($\tau$), depending on the specific implementation of the quantile loss function.
@@
To train the model, we minimise the loss function using stochastic gradient descent with mini-batches of size $B=8192$ observations. Specifically, we use the Adam optimiser (Kingma and Ba (2014)) with an initial learning rate of $0.0003$ and $(\beta_{1},\beta_{2})=(0.9,0.999)$. We adjust the number of epochs based on the number of observations available at each training period. The number of epochs is set to $100*(A/n)$, where $n$ is the number of observations available for training and $A$ is the adjustment constant equal to $3000000$ for the full sample and $1500000$ for the fluid sample, derived from the average number of observations in each sample over time. [^7] See Section 3.1 for the description of the data used. We use early stopping with a patience of 2 epochs, where 20% of the training data is used as validation to determine the stopping point. After early stopping, the model state with the lowest validation loss is taken as the final model. We use Leaky Relu (Maas et al. (2013)) as the activation function in all layers except the last layer in each subnetwork. We use batch normalisation (Ioffe and Szegedy (2015)) in all layers except the last layer in both the standardised $\tau$-quantiles subnetwork and the market volatility subnetwork. We apply dropout to all layers except the last layer in each subnetwork and use a rate of $0.2$ as a result of the hyperparameter search described in Appendix D. We also apply an L1 penalty to the weights of the first stage in each subnet. The L1 penalty for the first layer is $0.0001$, and for the second layer, it is $0.00001$. The first layer of the market volatility network is also regularised by an L2 penalty of $0.00001$. As an additional form of regularisation, we use an ensemble of 20 networks.
@@
The model is trained from scratch for the first training period and then fine-tuned for subsequent training periods [^8] Initial weights are taken from the previous training period.. Fine-tuning greatly reduces the computational burden of having to re-estimate the model each period and is a natural choice given the large overlap of subsequent training samples with the widening estimation window approach. It is expected that one year of new data should not change the estimated parameters too much unless there is a novel high-information event such as the dot-com bubble or the global financial crisis, in which case fine-tuning requires more epochs to retrain.
@@ 2.3 Approximating probability density and its moments from quantiles using B-Splines
To approximate the probability density, let’s consider a grid of $\tau$ quantile forecasts $\{Q_{j}(\tau),\tau_{j}\}_{j=1}^{K}$, where $\tau_{j}$ is the probability and $Q_{j}(\tau)$ is the corresponding quantile value, containing $K$ grid points that can be used to obtain the cumulative distribution function $F:\mathbb{R}\rightarrow[0,1]$. Since stock returns range from $-1$ to $\infty$, from a practical perspective we are interested in $F:[-1,\infty)\rightarrow[0,1]$. To get a denser representation, we use cubic B-splines [^9] We use the splrep and BSpline functions from the Python scipy library, Virtanen et al. (2020). to interpolate between adjacent quantiles by constructing piecewise polynomial functions within each interval $[Q_{j},Q_{j+1}]$. Instead of interpolating the predicted quantiles directly, we create a new denser $\mathcal{T}^{d}$ grid with one hundred equally spaced points between every two original quantile prediction points, except the largest and smallest, i.e. the denser grid has $(K-3)\times 100$ values, and fit the cubic B-spline to approximate the cumulative distribution function. To derive the probability density function, the derivative of the spline for the base function is calculated at each interval to obtain the new spline function representing the probability density function.
Figure 3 shows an example of the interpolated cumulative distribution function on the left and the implied probability density function on the right. The figure is based on $\tau$-quantile forecasts of 22-day (monthly) returns for Microsoft in October 2008. The density function is relatively close to a Student’s t-distribution with noticeably large tails during the depths of the Great Financial Crisis.
@@
As the interpolated density function does not cover the extreme tails of the distribution beyond the smallest and largest predicted quantile, these tails are modelled by discrete probabilities for values at the ends of the new $\mathcal{T}^{d}$ grid. We use discrete probabilities implied by the estimated spline function. The density function is truncated for returns less than -1, and discrete probability is then considered for $-1$ returns. Excluding the extreme tails from the density function results in a slight underestimation of the variance and a more pronounced underestimation of the kurtosis of the distribution. Inferring the distribution in the tails from empirical data is problematic, so we chose this solution to avoid the problem to some extent. In cases where B-spline interpolation does not produce plausible results, we use a linear B-spline function as a fallback (see Appendix E.1 for details). A complete algorithm for approximating the probability density function from the quantile forecasts is described in detail in Appendix E, Algorithm 1.
Importantly, the B-spline approximation allows us to obtain moments of the probability distribution function. We first derive the $z$th non-central moment of a density function $f(x)$

### v2

@@ 2.1 Network architecture
While neural networks have been successfully used in stock return forecasting (Gu et al., 2020a), quantile forecasting is challenging because we need to capture both cross-sectional variation across stocks and market-wide variation over time. One way to address this challenge is to use cross-sectional standardisation of returns to remove the time series noise. This is possible from an asset pricing perspective, where the goal is to rank individual stocks cross-sectionally to create long-short portfolios. However, since our goal is to forecast quantiles of future raw (non-standardised) stock returns, we propose a two-stage neural network architecture that exploits the predictability of standardised returns in the first stage and generates quantile forecasts of true (non-standardised) stock returns in the second stage. In this way, we allow the neural networks to exploit the information in standardised returns and to adjust for expected market volatility.
@@
where $\overline{\sigma}_{t}=\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\sigma_{i,{t}}$ is a cross-sectional average of stock-level volatilities, $\sigma_{i,t}$ is a stock-level volatility estimate of stock $i$ at period $t$, and $N_{t}$ is the number of stocks in the panel at period $t$. Ideally, we would use $\overline{\sigma}_{t+1}$ instead of $\overline{\sigma}_{t}$ to rescale the standardised quantile forecasts back to the raw quantile forecasts of stock returns, but this is not known at period $t$ when the forecast is made. Therefore, we use the cross-sectional average of stock-level volatilities at period $t$ instead of the future value of $\overline{\sigma}_{t+1}$ to avoid look-ahead bias. $\sigma_{i,t}$ is estimated using an exponentially weighted moving average of the squared returns of stock $i$ with a decay factor of 0.940.940.94 and is one of the input variables to the model. [^3] The 0.94 parameter is widely used in both academia and the financial industry, an example being the RiskMetrics GARCH specification. See Appendix J for a visualisation of $\overline{\sigma}$ over time in the US and more detailed arguments for the two-stage setup.
The idea behind standardising individual stock returns using the cross-sectional average of stock-level volatilities rather than the stock volatility itself is to reduce the noise from individual stock volatility estimates. For the same noise-reducing reason, we standardise stock returns only by dividing by the average volatility and not by subtracting the average return, either stock or market, because average returns are not persistent.
Let $\mathcal{T}=\{\tau_{1},\ldots,\tau_{K}\}$ be a set of quantile levels with $\tau_{k}\in(0,1)$. A stock return quantile is then modelled by the two-stage quantile neural network with two sub-networks as
$Q_{r_{i,t+1}}^{\mathcal{T}}$ $=$ $\Big{\{}Q_{r_{i,t+1}}(\tau_{1}|x_{i,t},z_{t},\overline{\sigma}_{t}),\ldots,Q_{r_{i,t+1}}(\tau_{K}|x_{i,t},z_{t},\overline{\sigma}_{t})\Big{\}}$ (2)
$=$ $\Bigg{[}\Big{(}\underbrace{g^{(L)}_{W_{S_{1}}^{(L)},b_{S_{1}}^{(L)}}\circ\ldots\circ g^{(1)}_{W_{S_{1}}^{(1)},b_{S_{1}}^{(1)}}\left(x_{t}\right)}_{\text{standardised $\tau$ quantiles sub-network}}\Big{)}\times\overline{\sigma}_{t}\Bigg{]}\times\Big{(}\underbrace{g^{(2)}_{W_{S_{2}}^{(2)},b_{S_{2}}^{(2)}}\circ g^{(1)}_{W_{S_{2}}^{(1)},b_{S_{2}}^{(1)}}\left(z_{t}\right)}_{\text{market-wide volatility sub-network}}\Big{)}$
$=$ [$\underbrace{\Bigg{[}\Big{\{}\underbrace{Q_{\widetilde{r}_{i,t+1}}(\tau_{1}|x_{i,t}),\ldots,Q_{\widetilde{r}_{i,t+1}}(\tau_{K}|x_{i,t})}_{\text{Stage I}}\Big{\}}\times\overline{\sigma}_{t}\Bigg{]}\times\widehat{\sigma}_{t}^{M}}_{\text{Stage II}},$
@@
It is important to note that, in contrast to the literature, we consider a multi-output (deep) neural network to characterise the collection of quantiles. Our network has two sub-networks, namely a standardised $\tau$ quantile network and a market-wide volatility network [^4] Note that the market-wide volatility sub-network is forecasting optimal multiplicative adjustment factor to $\overline{\sigma}_{t}$ to produce quantile predictions for non-standardised stock returns. It does not forecast market volatility directly. in which $l\in{1,...,L}$ hidden layers transform the input data into a chain using a collection of nonlinear activation functions $g_{S}^{(1)},\ldots,g_{S}^{(L)}$. An activation function, $g^{(\ell)}_{W_{S}^{(\ell)},b_{S}^{(\ell)}}$, is used as a
@@
$\overline{\sigma}_{t}$ $Q^{\mathcal{T}}_{\widetilde{r}_{i,t}}$ (37$\times$1) $\times$$\times$$\times$(128$\times$2)(128$\times$2) Bottleneck (4$\times$1) $x_{t}$ (176$\times$1) $\times$$\times$$\times$ $z_{t}$ (18$\times$1) (8$\times$1) $\widehat{\sigma}_{t}^{M}$ (1$\times$1) $Q^{\mathcal{T}}_{r_{i,t}}$ (37$\times$1) Stage IStage II Standardized $\tau$-quantiles sub-network Market-wide volatility sub-network
[…]
Figure 1 details the architecture of the two-stage quantile neural network model defined above. This is our main model used for quantile forecasting, and we refer to it as the two-stage model throughout the paper. The first stage of the model is used to forecast quantiles of cross-sectionally standardised next month stock returns $Q^{\mathcal{T}}_{\widetilde{r}_{i,t}}$. In our empirical application, we use 37 $\tau$s ranging from the extreme left tail of the distribution (0.00005) to the extreme right tail (0.99995).
Once the set of quantile levels $\mathcal{T}$ is specified, the two-stage architecture proceeds to generate forecasts for each quantile through a sequence of rescaling operations. The second stage of the model generates quantile forecasts of the next month’s raw stock returns $Q_{r_{i,t}}(\tau)$ for $\tau\in\mathcal{T}$. This stage consists of three steps. First, the standardised quantile forecasts $Q^{\mathcal{T}}_{\widetilde{r}_{i,t}}$ from the first stage and the cross-sectional average of stock-level volatilities $\overline{\sigma}_{t}$ are used as inputs to rescale the standardised quantile forecasts back to the raw stock return scale $Q_{r_{i,t}}(\tau)=Q_{\widetilde{r}_{i,t}}(\tau)\cdot\overline{\sigma}_{t}$. Second, the market-wide volatility sub-network generates a second multiplicative scaling factor to reintroduce a forward-looking aspect to the market-wide volatility scalar $\overline{\sigma}_{t}$. This market-wide volatility scaling factor is the same for all stocks in a given period $t$ and is used to refine the quantile forecasts rescaled in the first step. Finally, both the standardised and raw quantile forecasts are clipped at -100% during the forward pass of the neural network.
@@ 2.2 Training with quantile loss function
In order to jointly predict multiple $\tau$-quantiles using a two-stage neural network, individual $\tau$-quantile losses must be aggregated into a single loss. Our model outputs both standardised and raw stock return quantiles, and we combine the respective losses using a simple average. Let $\mathcal{T}=\{\tau_{1},\ldots,\tau_{K}\}$ denote a set of $K$ quantile levels, where each $\tau_{k}\in(0,1)$.
@@
Here, $B$ is the batch size, $K$ the number of quantiles in $\mathcal{T}$, and $\rho_{\tau}(\cdot)$ is the quantile loss function:
$\rho_{\tau}(\xi)=\mathds{1}_{(\xi\geq 0)}\tau\xi+\mathds{1}_{(\xi<0)}(\tau-1)\xi$ (4)
This piecewise linear function penalises underprediction and overprediction asymmetrically based on the quantile level $\tau$, with equal weighting across stocks and quantiles in the aggregate loss. [^5] The linear combination of losses from the two output heads uses equal weights, though in practice, the standardised return component dominates due to differences in scale. This design choice aligns with our modelling rationale, where the first stage captures most of the learning and the second stage refines the result. Although $\rho_{\tau}(\cdot)$ is not differentiable at zero, modern autodiff frameworks such as PyTorch handle subgradients at non-differentiable points automatically.
@@
To train the model, we minimise the loss function using stochastic gradient descent with mini-batches of size $B=8192$ observations. Specifically, we use the Adam optimiser (Kingma and Ba (2014)) with an initial learning rate of 0.00030.00030.0003 and $(\beta_{1},\beta_{2})=(0.9,0.999)$. We adjust the number of epochs based on the number of observations available at each training period. The number of epochs is set to $100*(A/n)$, where $n$ is the number of observations available for training and $A$ is the adjustment constant equal to 300000030000003000000 for the full sample and 150000015000001500000 for the liquid sample, derived from the average number of observations in each sample over time. [^6] See Section 3.1 for the description of the data used. We use early stopping with a patience of 2 epochs, where 20% of the training data is used as validation to determine the stopping point. After early stopping, the model state with the lowest validation loss is taken as the final model. We use Leaky Relu (Maas et al. (2013)) as the activation function in all layers except the last layer in each sub-network. We use batch normalisation (Ioffe and Szegedy (2015)) in all layers except the last layer in both the standardised $\tau$-quantiles sub-network and the market-wide volatility sub-network. We apply dropout to all layers except the last layer in each sub-network and use a rate of 0.20.20.2 as a result of the hyperparameter search described in Appendix D. We also apply an L1 penalty to the weights of the first stage in each subnet. The L1 penalty for the first layer is 0.00010.00010.0001, and for the second layer, it is 0.000010.000010.00001. The first layer of the market-wide volatility network is also regularised by an L2 penalty of 0.000010.000010.00001. As an additional form of regularisation, we use an ensemble of 20 networks.
@@
The model is trained from scratch for the first training period and then fine-tuned for subsequent training periods, meaning that weights from the previous training period are used to initialise the network instead of reinitialising randomly. Fine-tuning greatly reduces the computational burden of having to re-estimate the model each period and is a natural choice given the large overlap of subsequent training samples with the widening estimation window approach. It is expected that one year of new data should not change the estimated parameters too much unless there is a novel high-information event such as the dot-com bubble or the global financial crisis, in which case fine-tuning requires more epochs to retrain.
@@ 2.3 Approximating probability density and its moments from quantiles using B-Splines
To approximate the probability density, let’s consider a grid of $\tau$ quantile forecasts $\{Q_{j}(\tau),\tau_{j}\}_{j=1}^{K}$, where $\tau_{j}$ is the probability and $Q_{j}(\tau)$ is the corresponding quantile value, containing $K$ grid points that can be used to obtain the cumulative distribution function $F:\mathbb{R}\rightarrow[0,1]$. Since stock returns range from 1-1- 1 to $\infty$, from a practical perspective we are interested in $F:[-1,\infty)\rightarrow[0,1]$. To get a denser representation, we use cubic B-splines [^7] We use the splrep and BSpline functions from the Python scipy library, Virtanen et al. (2020). to interpolate between adjacent quantiles by constructing piecewise polynomial functions within each interval $[Q_{j},Q_{j+1}]$. Instead of interpolating the predicted quantiles directly, we create a new denser $\mathcal{T}^{d}$ grid with one hundred equally spaced points between every two original quantile prediction points, except the largest and smallest, i.e. the denser grid has $(K-3)\times 100$ values, and fit the cubic B-spline to approximate the cumulative distribution function. To derive the probability density function, the derivative of the spline for the base function is calculated at each interval to obtain the new spline function representing the probability density function.
@@
Figure 2 shows an example of the interpolated cumulative distribution function on the left and the implied probability density function on the right. The figure is based on $\tau$-quantile forecasts of 22-day (monthly) returns for Microsoft in October 2008. The density function is relatively close to a Student’s t-distribution with noticeably large tails during the depths of the Great Financial Crisis. Figure 3 then complements this with the evolution of the forecasted distribution of the 22-day-ahead returns over the period 1996-2020.
Figure 3: Example of forecasted probability density function over time. This figure displays the forecasted probability density function for out-of-sample 22 days-ahead returns of an Microsoft stock (CUSIP 59491810) generated using our two-stage quantile neural network model.
As the interpolated density function does not cover the extreme tails of the distribution beyond the smallest and largest predicted quantile, these tails are modelled by discrete probabilities for values at the ends of the new $\mathcal{T}^{d}$ grid. We use discrete probabilities implied by the estimated spline function. The density function is truncated for returns less than -1, and discrete probability is then considered for 1-1- 1 returns. Excluding the extreme tails from the density function results in a slight underestimation of the variance and a more pronounced underestimation of the kurtosis of the distribution. Inferring the distribution in the tails from empirical data is problematic, so we chose this solution to avoid the problem to some extent. In cases where B-spline interpolation does not produce plausible results, we use a linear B-spline function as a fallback (see Appendix E.1 for details). A complete algorithm for approximating the probability density function from the quantile forecasts is described in detail in Appendix E, Algorithm 1.
Importantly, the B-spline approximation allows us to obtain moments of the cumulative probability distribution function. We first derive the $z$th non-central moment of a density function $f(x)$